What is the resulting density when you mix 5.0 L of ethyl alcohol and 12.0 L of water?


1
Expert's answer
2021-03-03T11:25:28-0500

Density of ethyl alcohol = 0.789 g/mL

Density of water = 1.0 g/mL

"V_{e.a.} = 5000 \;mL \\

V_w = 12000 \;mL \\

V_{total} = 17000 \;mL \\

m_{e.a.} = 0.789 \times 5000 = 3945 \;g \\

m_w = 1.0 \times 12000 = 12000\; g \\

m_{total} = 3945 + 12000 = 15945 \;g"

The resulting density "= \frac{15945}{17000} = 0.938 \;g/mL"

Answer: 0.938 g/mL